DIRC Operations Report
Wednesday 2008/01/30 to Wednesday 2008/02/06
Oncall: Jennifer Prendki (Joe Schwiening on backup)
Fast Monitoring: Joe Schwiening and Jennifer Prendki
Luminosities this week
PEP peak lumi: ~11.0 x 1033
(last week: ~9.5 x 1033)
PEP delivery: 4.5 fb-1
BaBar integrated lumi: 4.21 fb-1
Run 7 total delivery: 22.17 fb-1
Run 7 total BaBar recorded: 21.34 fb-1 total,
including 1.80 fb-1 off-peak data and including 18.92
fb-1 on the Y(3S)
PEP activities/issues this week
- Vacuum leak problems in PEP have been bothering all week
long. They were fixed yesterday.
BaBar activities/issues this week
- Some Y(3S) offpeak data has been taken last week-end. The
data quality of those runs looks fine.
DIRC activities/issues this week
See DIRC logbook for details
Lumi and Background this week
- Joe and Jennifer rotating oncall and backup duties. Joe is now on-call, Jennifer in back-up.